OpenStudy (anonymous):

A bank pays interest at the nominal rate of 3.8% per year. What is the effective annual yield if compounding is: annual? monthly? continuously?

OpenStudy (anonymous):

monthly

OpenStudy (anonymous):

no, I'm suppose to say what the yield is for each one. whats the yield if its compounded annualy, monthly and continously

OpenStudy (anonymous):

its monthly

ganeshie8 (ganeshie8):

Annual yield = \(\large \left( 1+ \dfrac{r}{n} \right)^n - 1\)

ganeshie8 (ganeshie8):

\(r = 0.038\) fixed quantity : \(\large \left( 1+ \dfrac{0.038}{n} \right)^n - 1\)

ganeshie8 (ganeshie8):

change \(n\) based on compounding and evaluate

ganeshie8 (ganeshie8):

\(n\) is the compounding interval, for annual compounding, \(n = 1\)

OpenStudy (anonymous):

how would you do continuous?

ganeshie8 (ganeshie8):

for continuous use the method we have used in ur previous problem

OpenStudy (anonymous):

i have 3 min to submit these can you solve them out? I keep getting them wrong

ganeshie8 (ganeshie8):

yield = \(\large e^{r} - 1\)

ganeshie8 (ganeshie8):

\(\large e^{0.038}-1\)

ganeshie8 (ganeshie8):

feed it to wolfram : http://www.wolframalpha.com/input/?i=+e%5E%7B0.038%7D-1

ganeshie8 (ganeshie8):

maybe they want answers in percentages... so multiply them by 100

OpenStudy (anonymous):

your answers were right, unfortunately they didn't make it in on time. but I got everything else right. Thanks for trying though. I hate how I never have enough time to do these :/

ganeshie8 (ganeshie8):

okk